OpenStudy (anonymous):

What is the simplified version of 3 log x + log y + log z ? help pleasee

OpenStudy (saranya):

log (x^m)=mlogx log(m)+log(n)=log(mn) log(m)-log(n)=log (m/n) 3logx=log(x^3) and -logy-logz=-(logy+logz)=-log(yz) 3logx-logy-logz=log(x^3)-log(yz) =log(x^3/yz)

ganeshie8 (ganeshie8):

use the logarithms property : log m + log n = log(mn)

OpenStudy (anonymous):

thankss